The sum of two numbers is 93. The second is 3 more than 4 times the first. What are the two numbers?
Answer by BrittanyM(80)

We can set up a system of equations to find the vaulues of the two unknown numbers.

The sum of two numbers is 93:
x + y = 93
The second is three more than four times the first;
y = 4x + 3

We can substitute y in the first equation in order to find the exact value of x.
x + (4x + 3) = 93
5x + 3 = 93
5x = 90
x = 18

Now we can plug this into the equation to find the exact value of y.

(18) + y = 93
y = 75